Solve:(5w + 2)^2 - 3(5w + 2) - 4 = 0
Appreciate all help/feedback.

Answer by mananth(16946)

(5w + 2)^2 - 3(5w + 2) - 4 = 0

let (5w+2)=a for simplification
a^2-3a-4
a^2-4a+a-4 =0
a(a-4)+1(a-4)=0
(a-4)(a+1)=0
(5w+2-4)(5w+2+1)=0
(5w-2)(5w+3)=0
w=2/5 OR w=-3/5